OpenStudy (anonymous):

Finding electric potential?? There are two thin, conducting, spherical shells. The inner shell with radius 15cm has a charge of 10 nC and the outer one has r=30cm and charge of 15 nC. How do you find the e. potential in each region?

OpenStudy (anonymous):

How do you find the electric potential in the region in between the smaller and the larger shells?

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

I have the correct electric fields for each region: A= 0 v/m B= 90/(r^2) V/m C= -45/(r^2) V/m